Circular No. 7593

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001aa IN UGC 10888 G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong of an apparent supernova (mag ∼ 18.0) on seven unfiltered CCD frames taken from Mar. 5.124 to 5.222 UT. Confirming images were obtained by Armstrong in poor conditions on Mar. 6.1, showing the new object at mag 18.4. SN 2001aa is located at α = 17^{h}29^{m}57^{s}.71, δ = +60^{o}21′23′′.2 (equinox 2000.0; mean of four measurements), which is 10′′ west and 22′′ north of the center of UGC 10888. Nothing is visible at this location on earlier CCD images taken by Armstrong on Jan. 13 (limiting mag 19.0) and 16 (limiting mag 19.5), neither is anything visible on Palomar Sky Survey images taken on 1987 July 27 (limiting blue mag 22.5), 1992 May 10 (red mag 20.8), and 1992 June 20 (infrared mag 19.5). A red star (mag 18.8) located 9′′ north of SN 2001aa is listed in the USNO A2.0 catalogue.

SUPERNOVA 2001X IN NGC 5921 R. Chornock, W. D. Li, and A. V. Filippenko, University of California

at Berkeley, report that a spectrum of SN 2001X (cf. IAUC 7591), obtained with the Lick Observatory Shane 3-m telescope on Mar. 3.5 UT under poor conditions, shows a very blue continuum with strong H Balmer lines, indicating that it is a type-II supernova before maximum light. The He I 587.6-nm line is also strong.

SUPERNOVA 2001Y IN NGC 3362 M. Phillips, G. Martin, and G. Valladares, Carnegie Observatories, ob-

tained a spectrum (range 350–850 nm) of SN 2001Y (cf. IAUC 7592) on Mar. 5.35 UT with the 6.5-m Baade telescope at Las Campanas Observatory. The spectrum, which reveals broad H I and Fe II emission lines on a strong, blue continuum, closely resembles that of the peculiar type-II

event SN 1999gk (see http://cfa-www.harvard.edu/cfa/oir/Research/
supernova/spectra/sn99gk.gif).

SUPERNOVA 2001Z IN IC 3528 Phillips, Martin, and Valladares also obtained a spectrum of SN 2001Z

(cf. IAUC 7592) on Mar. 5.37 UT with the Baade telescope. This shows weak, relatively narrow Hα and Hβ emission lines on a strong, blue continuum, consistent with a type-II supernova caught at a very early epoch. The redshift of the emission lines is z = 0.045.

2001 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green